# White-Basilisk: A Hybrid Model for Code Vulnerability Detection

### v1

@@ 5 Experimental Results: Small Model, Big Impact
White-Basilisk 0.9630 0.2907 0.7239 0.9942 0.9490 0.0398
CodeT5 0.9667 0.1970 0.8993 0.9567 0.6493 0.7730
CodeBERT 0.9687 0.2086 0.8878 0.9557 0.6288 0.8177
UnixCoder 0.9686 0.2143 0.8921 0.9646 0.6546 0.6230
UnixCoder w/ balancing 0.9599 0.2628 0.8849 - - -
StarCoder2 0.9702 0.1805 0.8964 0.9620 0.6826 0.6914
CodeGen2.5 0.9665 0.1961 0.9151 0.9657 0.6730 0.6173
LineVul - - - - 0.9100 0.1400
Table 2: Draper, REVEAL, and VulDeepecker Evaluation Results
@@
F1 Acc F1 F1 Prec
White-Basilisk 0.6069 0.8988 0.4934 0.9388 0.9720
Russell et al. (2018) 0.5660 - - - -
VulBERTa-MLP 0.4334 0.8448 0.4527 0.9303 0.9576
VulBERTa-CNN 0.5792 0.7973 0.4259 0.9086 0.9526
Baseline-BiLSTM 0.4684 0.7713 0.3911 0.6697 0.5258
Baseline-TextCNN 0.4940 0.7322 0.3741 0.7580 0.6348
REVEAL - 0.8437 0.4125 - -
VulDeepecker - - - 0.9290 0.9190
5.1 Ablation Studies: Validating Key Claims
To further validate our architectural choices and claims, we conducted three targeted ablation studies. Detailed results and analysis for each study are available in the Appendixes H, I, J.
5.1.1 Unified Model Analysis: Can One Model Rule Them All?
We trained a unified White-Basilisk model on the combined training data from all five datasets to assess cross-dataset generalization. This model maintained robust performance across diverse vulnerability types, with F1 scores of 0.938 on BigVul, 0.925 on VulDeepecker, 0.549 on Draper, 0.470 on REVEAL, and 0.233 on PRIMEVUL. Notably, the model maintained near-optimal performance on larger datasets while showing reasonable transfer to more challenging ones, demonstrating White-Basilisk’s ability to learn generalizable vulnerability patterns despite its compact size. The model achieved perfect detection (F1=1.000) for 22 different CWE categories in BigVul, including critical issues like out-of-bounds write (CWE-787) and cryptographic vulnerabilities (CWE-310).
Table 3: Performance on Key CWEs Across Datasets
CWE Description Datasets Samples F1 Range
CWE-119 Buffer Overflow BigVul, VulDeep, Draper 15,584 0.748-0.969
CWE-399 Resource Management BigVul, VulDeep 7,031 0.780-0.923
CWE-264 Access Control BigVul 1,240 0.925
CWE-476 NULL Pointer Deref. BigVul, Draper 1,709 0.686-0.971
CWE-416 Use After Free BigVul 963 0.958
5.1.2 Attention Efficiency Analysis: Linear vs. Quadratic Scaling
Does White-Basilisk’s linear-complexity attention truly enable efficient long sequence processing without performance degradation? We compared our approach against standard quadratic attention, measuring both memory usage and predictive performance across varying sequence lengths. As shown in Table 4, our approach achieves dramatic memory savings—using 24× less memory (1,338MB vs. 32,409MB) for sequences under 16K tokens while maintaining comparable or superior F1 scores (0.943 vs. 0.937). Crucially, while standard attention fails with out-of-memory errors beyond 16K tokens, White-Basilisk successfully processes sequences up to 131K tokens with only modest memory growth, demonstrating true linear scaling.
Most impressively, performance analysis revealed that White-Basilisk maintains strong detection capabilities even for extremely long sequences, with F1 scores of 0.943 for sequences under 16K tokens, 0.800 for 16K-32K tokens, and a perfect 1.000 for 32K-65K tokens. This confirms our architecture’s ability to capture long-range dependencies crucial for identifying vulnerabilities that span multiple functions or files. For detailed sequence-length analysis across all datasets, see Appendix I.
Table 4: Memory Usage and Performance by Sequence Length and Attention Type
Length Peak Memory (MB) BigVul F1 Score
Ours Standard Ours Standard
0-16K 1,338 32,409 0.943 0.937
16K-32K 1,654 OOM 0.800 -
32K-65K 2,213 OOM 1.000 -
65K-131K 3,322 OOM - -
OOM: Out of Memory on NVIDIA A100 40GB GPU
5.2 Efficiency and Environmental Impact
The remarkable performance of White-Basilisk becomes even more significant when considering its computational efficiency. At approximately 200M parameters, our model is orders of magnitude smaller than models like StarCoder2 (7B) while achieving superior results. This efficiency translates to substantial environmental benefits, as shown in Table 5.
Table 5: Comparison of CO2 Emissions
White-Basilisk CodeBERT StarCoder UnixCoder CodeT5 Gasoline Car (per Year) CO2 (kg) 85.5 2,240 23,000,000 2,048 1,136 4,600
We estimated the CO2 emissions during training using the Machine Learning Impact calculator (Lacoste et al., 2019). The environmental impact of White-Basilisk (85.5kg CO2) represents just 3.8% of CodeBERT’s emissions (2,240kg) and a mere 0.0004% of StarCoder’s emissions (23,000,000kg). This stark contrast underscores the environmental benefits of our approach, with CO2 savings equivalent to taking hundreds of cars off the road compared to larger model training.
The efficiency gains extend to inference as well. White-Basilisk can process sequences up to 128K tokens on a single NVIDIA A100 GPU, enabling whole-codebase analysis without the distributed computing infrastructure required by larger models. This practical efficiency makes advanced vulnerability detection accessible to a broader range of organizations and researchers.
Our results challenge the prevailing "bigger is better" paradigm in AI model development. White-Basilisk demonstrates that carefully designed architecture, rather than simply increasing parameter count, can yield superior performance while drastically reducing computational and environmental costs. This finding has significant implications for sustainable AI development, suggesting that the path forward may lie in architectural innovation rather than ever-larger models.
@@ 6 Limitations and Future Work
While White-Basilisk shows promising results in code vulnerability detection, it is important to acknowledge its current limitations and outline future directions. The main limitation of White-Basilisk is its focus on C and C++ codebases. The model’s ability to generalize across a broader range of programming languages, especially those with different syntaxes or paradigms, warrants further exploration. This constraint, combined with potential biases in our training and evaluation datasets, may limit the model’s generalizability to diverse real-world codebases. To address this, future work will involve expanding the model’s training to include a wider variety of programming languages and curating more representative datasets that reflect a broader spectrum of code samples and vulnerability types.
Also, despite strong performance metrics, White-Basilisk is not infallible. False positives and false negatives, particularly when detecting novel or zero-day vulnerabilities, remain an ongoing challenge. Additionally, while the model is capable of processing long sequences, its true understanding of complex, long-range dependencies in code still needs further investigation. Our future research will focus on reducing error rates, especially in high-stakes scenarios, and enhancing the model’s ability to analyze convoluted code structures spanning multiple functions or files.
Another area for improvement is the model’s explainability. Currently, White-Basilisk’s decision-making process is not sufficiently transparent. Improving the model’s ability to provide clear, actionable explanations for detected vulnerabilities is essential for building trust and delivering meaningful insights to developers. Future work will explore methods to offer context-aware understanding of the potential impact of vulnerabilities, as well as suggested fixes, ultimately aiming to evolve White-Basilisk into a comprehensive code analysis assistant rather than a mere detection tool.
White-Basilisk’s performance on a relatively small training dataset (2M samples) is impressive, but raises questions about potential limitations in its knowledge base compared to models trained on much larger datasets. To address this, we plan to scale White-Basilisk to approximately 1 billion parameters. While still modest in comparison to larger language models, this increase in parameter count aims to significantly boost performance while continuing to challenge the notion that only the largest models can deliver state-of-the-art results.
Beyond code vulnerability detection, there is potential for White-Basilisk’s architecture and training approach in broader AI applications. Future research will investigate its efficacy in various NLP tasks, exploring whether its computational efficiency and long-context capabilities can offer more resource-efficient alternatives to existing LLMs.
@@ 7 Conclusion
White-Basilisk represents a significant advancement in efficient AI model design, demonstrating that carefully architected compact models can outperform much larger alternatives in specialized domains. By achieving state-of-the-art results in code vulnerability detection with only 200M parameters, our work challenges the prevailing "bigger is better" paradigm in AI development. The model’s impressive performance—particularly its ability to process sequences up to 128,000 tokens—proves that context handling limitations can be addressed through architectural innovation rather than simply increasing model size.
The model’s extended context window addresses a fundamental limitation of many current LLMs, which often struggle with long-range dependencies and global code structure understanding. By enabling the analysis of entire codebases in a single pass, White-Basilisk can capture complex interdependencies and identify vulnerabilities that span multiple functions or files, a capability that has long eluded traditional approaches.
While the context-handling capabilities of White-Basilisk are its standout feature, it’s worth noting that these achievements have been realized with a relatively compact model of 200M parameters. This efficiency demonstrates that advances in AI are not solely dependent on increasing model size, but can also stem from innovative architecture design and training methodologies.
The implications of White-Basilisk’s approach extend beyond code vulnerability detection. The ability to handle extended contexts efficiently could prove valuable in numerous domains where long-range understanding is crucial, such as document analysis, complex system modeling, or long-form text generation. Moreover, the model’s efficiency opens up possibilities for deployment in resource-constrained environments, potentially bringing advanced AI capabilities to a broader range of applications and users.
In conclusion, White-Basilisk represents a significant step forward in addressing the context limitations of current LLMs, while also demonstrating that such advances need not come at the cost of excessive model size or computational requirements. As we continue to refine and expand upon this approach, we anticipate exciting developments in the field of AI, particularly in tasks that require deep understanding of extended contexts. The potential implications of this research are substantial, and we look forward to seeing how these ideas evolve and find application in diverse areas of AI and beyond.
@@ Appendix A Appendix
Appendix B Evaluation Metrics
@@ B.0.4 F1 Score
[…]
B.0.5 Vulnerability Detection Score (VD-S)
@@
Each metric serves a specific purpose in evaluating different aspects of model performance, from general classification accuracy to specialized vulnerability detection capabilities. The combination of these metrics provides a comprehensive assessment of a model’s effectiveness in real-world security applications.
Appendix C Dataset Statistics Analysis

### v4

@@ 6.1 Main Results: Dataset-Specific and Unified Training
CodeT5 0.967 0.197 0.899 0.957 0.649 0.773
CodeBERT 0.969 0.209 0.888 0.956 0.629 0.818
UnixCoder 0.969 0.214 0.892 0.965 0.655 0.623
StarCoder2 0.970 0.181 0.896 0.962 0.683 0.691
CodeGen2.5 0.967 0.196 0.915 0.966 0.673 0.617
White-Basilisk (Dataset-Specific) 0.963 0.291 0.724 0.994 0.949 0.040
White-Basilisk (Unified) 0.952 0.233 0.767 0.993 0.938 0.062
Table 2: Performance on traditional vulnerability detection benchmarks
@@
F1 Acc F1 F1 Precision
Russell et al. (2018) 0.566 - - - -
VulBERTa-MLP 0.433 0.845 0.453 0.930 0.958
VulBERTa-CNN 0.579 0.797 0.426 0.909 0.953
Baseline-BiLSTM 0.468 0.771 0.391 0.670 0.526
Baseline-TextCNN 0.494 0.732 0.374 0.758 0.635
REVEAL - 0.844 0.413 - -
VulDeePecker - - - 0.929 0.919
White-Basilisk (Dataset-Specific) 0.607 0.899 0.493 0.939 0.972
White-Basilisk (Unified) 0.549 0.888 0.470 0.925 0.939
White-Basilisk consistently outperforms all baseline models across every vulnerability detection benchmark. On PRIMEVUL, our dataset-specific model achieves F1=0.291, representing a 35.5% improvement over the best baseline (UnixCoder: F1=0.214). On BigVul, White-Basilisk achieves F1=0.949, dramatically outperforming all competing approaches. The model demonstrates similar superiority on traditional benchmarks.
The unified model demonstrates remarkable cross-dataset generalization capability, maintaining competitive performance across all benchmarks despite being trained on combined data from diverse sources. Performance differences between dataset-specific and unified training are modest, suggesting effective knowledge transfer without significant performance degradation. This generalization spans datasets with different characteristics demonstrating the robustness of our approach.
Analysis of vulnerability types reveals that White-Basilisk demonstrates strong performance across different Common Weakness Enumeration (CWE) categories, with particularly robust detection of memory-related vulnerabilities including CWE-119 (buffer overflows) and CWE-476 (NULL pointer dereferences). The unified model achieves perfect F1=1.000 scores for 22 different CWE categories in BigVul evaluation, with consistently high performance on access control issues (CWE-264: F1=0.925) and input validation vulnerabilities (CWE-20: F1=0.933). Detailed CWE-specific performance analysis is provided in Appendix I.
6.2 PRIMEVUL Paired Evaluation
To address concerns about potential dataset artifacts and superficial pattern recognition, we evaluated White-Basilisk on the PRIMEVUL paired dataset, which consists of vulnerable code paired with its patched counterpart. This evaluation specifically tests the model’s ability to distinguish between vulnerable code and its security fix, often involving subtle changes that challenge pattern-matching approaches. The evaluation uses four metrics: P-C measures correct classification of both vulnerable and patched code in each pair, P-V indicates cases where both functions are incorrectly classified as vulnerable, P-B represents cases where both functions are incorrectly classified as benign, and P-R captures complete prediction reversals where vulnerable code is classified as safe and patched code as vulnerable.
Table 3: PRIMEVUL paired evaluation results
Model Method P-C ↑ P-V ↓ P-B ↓ P-R ↓
White-Basilisk (Ours) Fine-tuning 12.92 42.08 42.92 2.08
GPT-4 Chain-of-Thought 12.94 54.26 24.47 8.33
GPT-4 Two-shot 5.14 71.63 21.45 1.77
CodeGen2.5 Fine-tuning 3.01 10.82 84.22 1.95
StarCoder2-7B Fine-tuning 2.30 8.16 88.30 1.24
CodeBERT Fine-tuning 1.77 11.35 86.17 0.71
Random Baseline - 22.70 26.24 26.42 24.65
White-Basilisk achieves 12.92% P-C performance, representing a 4.29× improvement over the best open-source baseline (CodeGen2.5: 3.01%) and nearly matching GPT-4 with chain-of-thought reasoning (12.94%). The balanced error distribution between P-V (42.08%) and P-B (42.92%) indicates active vulnerability reasoning rather than biased default predictions. However, all models perform substantially below random baseline (22.70%), highlighting the fundamental difficulty of this task.
6.3 Context Length Impact
We conducted two experiments to validate our architectural choices and context processing claims. First, we evaluated our unified model across context lengths from 512 to 131K tokens on BigVul and PRIMEVUL datasets. Second, we performed full-file analysis on PRIMEVUL by cross-matching validation samples with their complete C/C++ source files and classifying entire files (up to 524K tokens) rather than isolated vulnerable functions.
Table 4: Performance across different context lengths and full-file validation
Context Length Baseline Models Context BigVul PRIMEVUL
Acc F1 Acc F1
512 CodeBERT, CodeT5 0.974 0.752 0.958 0.142
1,024 UnixCoder, VulBERTa 0.984 0.855 0.954 0.156
2,048 CodeGen2.5 0.990 0.909 0.953 0.205
4,096 StarCoder2 (sliding) 0.992 0.933 0.953 0.219
16,384 StarCoder2 (max) 0.993 0.942 0.952 0.235
32,768 - 131,072 White-Basilisk 0.993 0.943 0.952 0.233
Full-File Context Validation
524,288 PRIMEVUL Full Files - - 0.956 0.288
The context length experiment shows consistent performance improvements from 512 to 16K tokens (BigVul F1: 0.752→0.942; PRIMEVUL F1: 0.142→0.235), after which performance plateaus. This pattern aligns with our dataset analysis showing most samples fall within the 16K token range. The full-file experiment demonstrates a substantial F1 improvement (0.233→0.288) when analyzing complete source files instead of isolated functions.
These results validate two key aspects of our architecture. The linear scaling with context length confirms that vulnerabilities benefit from broader contextual information, while the dramatic improvement in full-file analysis indicates that many vulnerabilities manifest through inter-function dependencies invisible in isolated code fragments. The ability to process 524K-token sequences demonstrates the practical utility of our linear-scaling attention mechanism for complete codebase analysis, a capability that would be computationally infeasible with standard quadratic attention approaches. Additional sequence length analysis is provided in Appendix H.2.
@@ 7 Limitations and Future Work
While White-Basilisk shows promising results, several limitations constrain its current applicability. The primary limitation is its focus on C and C++ codebases, with unknown generalization to other programming languages with different syntaxes and vulnerability patterns. Combined with potential dataset biases, this may limit effectiveness on diverse real-world codebases.
Two fundamental dataset limitations constrain our evaluation. Current benchmarks contain primarily function-level code snippets rather than file-level samples, potentially missing crucial contextual information where vulnerabilities emerge from interactions between multiple functions or modules. Many vulnerabilities appear benign in isolation but become dangerous through complex inter-function dependencies. Additionally, existing datasets contain mostly sequences under 16K tokens, preventing full demonstration of White-Basilisk’s extended context capability and long-range dependency modeling. While our ablation studies show consistent improvements from 512 to 16K tokens, the scarcity of longer sequences means we cannot fully validate our architecture’s full potential. These represent broader field limitations rather than specific shortcomings of our approach.
Despite strong benchmark performance, White-Basilisk faces challenges common to automated vulnerability detection. False positives reduce developer trust while false negatives leave vulnerabilities undetected. Performance on novel vulnerability types or zero-day exploits remains uncertain. The model’s explainability is limited, it provides binary classifications without indicating reasoning, vulnerability locations, or risk assessments, creating barriers for security-critical deployment where developers need actionable insights.
We plan to address dataset limitations by developing comprehensive benchmarks with file-level samples preserving contextual information and longer sequences exercising our extended context capabilities. However, we first needed to establish performance on existing benchmarks for fair comparison with prior work. Future research will enhance explainability through attention visualization and program analysis integration, scale to approximately 1 billion parameters while maintaining efficiency, and investigate our hybrid architecture’s applicability to broader NLP tasks requiring efficient long-sequence processing.
@@ 8 Conclusion
White-Basilisk introduces a hybrid architecture that integrates Mamba layers, linear-scaling attention, and Mixture of Experts to achieve linear computational scaling with sequence length. This architectural design enables processing of extended sequences while maintaining compact parameter count, demonstrating that efficient architectural choices can overcome the quadratic complexity limitations of standard transformers.
The model achieves state-of-the-art performance across five vulnerability detection benchmarks, consistently outperforming substantially larger models. The significant performance gains from full-file context analysis validate that vulnerability detection benefits substantially from extended context processing capabilities that span multiple functions and broader code structure. These results establish that linear-complexity architectures can deliver superior performance on challenging downstream tasks requiring long-range dependencies. Code vulnerability detection represents a particularly demanding application where subtle patterns must be identified across extensive sequences, making it an effective testbed for evaluating architectural efficiency versus capability trade-offs.
The hybrid approach demonstrates broader applicability beyond security applications. Tasks requiring analysis of long documents, complex system modeling, or extended contextual understanding can benefit from similar architectural principles that maintain linear complexity with respect to sequence length while preserving representational capacity. The compact design enables deployment in resource-constrained environments without sacrificing performance, potentially expanding access to advanced sequence processing capabilities across diverse applications and computational budgets.
@@ Appendix A Appendix
Appendix B Architectural Details and Hyperparameters
For reproducibility purposes we provide all the model parameters and training hyperparameters in this section:
B.1 Classification Head
  1. 1.
  2. 2.
  3. 3.
  4. 4.
$x_{1}$ $=\text{Dropout}(\text{GELU}(W_{1}h+b_{1}))$ (7)
$x_{2}$ $=\text{Dropout}(\text{GELU}(W_{2}x_{1}+b_{2}))$
$x_{3}$ $=\text{LayerNorm}(x_{2})$
$y$ $=W_{3}x_{3}+b_{3}$
B.2 Hyperparameter Details
Table 5: Hyperparameters for the White-Basilisk Model
Hyperparameter Value
Vocabulary Size 65,540
Hidden Size 512
Intermediate Size 1,792
Number of Hidden Layers 12
Hidden Activation SiLU
Initializer Range 0.02
RMS Norm Epsilon $1\times 10^{-6}$
Attention Parameters
Attention Layer Offset 2
Attention Layer Period 8
Number of Attention Heads 16
Number of Key-Value Heads 8
Mixture of Experts (MoE) Parameters
Expert Layer Offset 1
Expert Layer Period 2
Number of Experts 8
Experts per Token 2
Router Auxiliary Loss Coefficient 0.001
Mamba Parameters
Mamba Convolution Dimension 4
Mamba State Dimension 16
Mamba Delta Rank 256
Mamba Expansion Factor 2
B.3 Pretraining Hyperparameters
B.4 Fine-tuning Hyperparameters
B.5 Handling Class Imbalance
$w_{c}=\frac{N}{2N_{c}}$ (8)
B.5.1 Weighted Random Sampling Implementation
We implement weighted random sampling using PyTorch’s WeightedRandomSampler WITHOUT replacement, which samples each example exactly once per epoch but reorders them based on importance weights. This approach provides two key benefit":
  1. 1.
    Temporal Concentration: Vulnerable samples appear earlier in the training epoch when gradients are typically more effective for learning
  2. 2.
    Batch Distribution Optimization: Maximizes the number of batches that contain vulnerable samples by preventing them from being wastefully clustered together in the same batches
To illustrate the effectiveness of this approach, consider the PRIMEVUL training dataset with 184,427 samples containing 3.02% vulnerable samples (5,569 vulnerable samples) and batch size 4:
  • Total batches per epoch: 46,107
  • Batches with at least vulnerable sample: 7.78% (3,585 batches)
  • Batches with 0 vulnerable samples: 92.22% (42,521 batches)
The weighted sampling with $w_{c}=\frac{N}{2N_{c}}$ gives vulnerable samples higher selection probability, ensuring maximum vulnerable-containing batches while front-loading their appearance in the epoch. While severe class imbalance means many batches will still contain only non-vulnerable samples, the weighted sampling optimizes the distribution of available vulnerable samples across the training process.
B.5.2 Weighted Loss Function
We implement a weighted loss function, modifying the standard cross-entropy loss by incorporating the class-specific weights $w_{c}$. This ensures that the loss contribution from vulnerable samples is appropriately scaled relative to their frequency in the dataset.
B.5.3 SIFT (Scale-Invariant Fine-Tuning)
We implement automated adversarial training using SIFT to improve the model’s resilience. SIFT operates by introducing small perturbations to the input during training, encouraging the model to learn more robust features. In our implementation, we added a PerturbationLayer into the model architecture, which applies learnable perturbations to the input embeddings. The training process was designed to minimize both the task loss and the adversarial loss, the latter being computed as the difference between predictions on clean and perturbed inputs.
Appendix C Evaluation Metrics
@@ C.1 Standard Classification Metrics
$\text{F1}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (12)
Vulnerability Detection Score (VD-S)
@@
C.2 Pairwise Evaluation Metrics
The pairwise evaluation metrics are specifically designed for vulnerability detection tasks involving pairs of vulnerable code and their corresponding patched versions. These metrics assess a model’s ability to distinguish between vulnerable code and its security fix, addressing concerns about superficial pattern recognition.
P-C (Pair-Correct)
P-C measures the percentage of code pairs where both the vulnerable and patched versions are correctly classified:
$\text{P-C}=\frac{\text{Pairs with both correct classifications}}{\text{Total pairs}}\times 100\%$ (14)
A higher P-C score indicates better overall discrimination capability between vulnerable and safe code.
P-V (Pair-Vulnerable)
P-V indicates the percentage of pairs where both vulnerable and patched code are incorrectly classified as vulnerable:
$\text{P-V}=\frac{\text{Pairs where both classified as vulnerable}}{\text{Total pairs}}\times 100\%$ (15)
This metric reveals tendency toward false positive predictions and failure to recognize security patches.
P-B (Pair-Benign)
P-B represents the percentage of pairs where both vulnerable and patched code are incorrectly classified as benign:
$\text{P-B}=\frac{\text{Pairs where both classified as benign}}{\text{Total pairs}}\times 100\%$ (16)
This metric indicates conservative prediction behavior and potential missed vulnerabilities.
P-R (Pair-Reversed)
P-R captures the percentage of complete prediction reversals where vulnerable code is classified as safe and patched code is classified as vulnerable:
$\text{P-R}=\frac{\text{Pairs with reversed classifications}}{\text{Total pairs}}\times 100\%$ (17)
This metric identifies the most problematic prediction pattern, suggesting fundamental misunderstanding of vulnerability characteristics.
Interpretation Guidelines
For pairwise metrics, the ideal performance would show high P-C values with low P-V, P-B, and P-R values. The distribution of error types (P-V vs. P-B vs. P-R) provides insights into model behavior: balanced P-V and P-B with low P-R suggests active reasoning rather than biased default predictions, while high P-R indicates serious model confusion about vulnerability patterns.
Each metric serves a specific purpose in evaluating different aspects of model performance, from general classification accuracy to specialized vulnerability detection capabilities and robustness against dataset artifacts. The combination of these metrics provides a comprehensive assessment of a model’s effectiveness in real-world security applications.
Appendix D Dataset Statistics Analysis